OpenStudy (anonymous):

Evaluate. 8!/5!3! 1 56 112

OpenStudy (anonymous):

?

OpenStudy (anonymous):

8*7*6*5*4*3*2*1 all over 5*4*3*2*1*3*2*1

OpenStudy (anonymous):

Then simplify.

OpenStudy (anonymous):

\[\frac{8\times 7\times 6\times 5!}{5!3!}\] is a start

OpenStudy (anonymous):

cancel \(5!\) and \(3!=6\) top and bottom and get \(8\times 7\)

OpenStudy (anonymous):

56